bemine11:

Solve -4b + 2 = 18. b =

simsharrison:

simple algebra

bemine11:

i am not good algebra

emilysmith1105:

ok first subtract 2 from both sides, it will leave you with -4b=16. then divide both sides by -4 and you get b= -16/4 or -4

simsharrison:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @bemine11 i am not good algebra \(\color{#0cbb34}{\text{End of Quote}}\) what is 4 times 4?

bemine11:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @simsharrison \(\color{#0cbb34}{\text{Originally Posted by}}\) @bemine11 i am not good algebra \(\color{#0cbb34}{\text{End of Quote}}\) what is 4 times 4? \(\color{#0cbb34}{\text{End of Quote}}\) 16

emilysmith1105:

so b= -4

simsharrison:

ok -4 times -4? = 16 so b=-4

lilaspie7:

yup

supie:

So the first step is to change the plus to a minus -4b - 2 = 18 subtract 2 from both sides of the equation, \(−4b+2−2=18−2\) Which equals −4b=16 So then divide both sides -4 \(\LARGE\frac{-4b}{-4}=\frac{16}{-4}\) \(\cancel{\color{red}{\LARGE\frac{-4b}{-4}}}=\frac{16}{-4}\) So \(\frac{-4b}{-4}\) cancels all you have left is \(\frac{16}{-4}\) simplify that then you have your answer.

bemine11:

thx @supie

supie:

np